A skier has a mass of 5200 grams is sliding down a snowy slope (a 13π/80 incline). (a) Determine the magnitude of the Forces acting on the skier in Newton. (b) Then determine the acceleration of the skier if there will be a wind pushing the skier down the slide with a Force of 6N. (c) the skier stops at the end of the incline, what is the minimum required force to move the skier. (assume coefficient of kinetic friction is 0.05, coefficient of static friction is 0.3 & gravity is 9.5m/s2 )


1
Expert's answer
2021-04-06T18:24:46-0400

a)


"W=mg=(5.2)(9.5)=49.4\ N\\N=mg\cos{\frac{13\pi}{80}}=(5.2)(9.5)\cos{\frac{13\pi}{80}}=43.1\ N\\F_f=\mu_kmg\cos{\frac{13\pi}{80}}\\=(0.05)(5.2)(9.5)\cos{\frac{13\pi}{80}}=2.2\ N"

b)


"5.2a=49.4\sin{\frac{13\pi}{80}}-2.2+6\\a=5.4\frac{m}{s^2}"

c)


"F=(0.3)(5.2)(9.5)\cos{\frac{13\pi}{80}}+49.4\sin{\frac{13\pi}{80}}+6\\F=43.1\ N"